# Transfer of quantum game strategies

### v1

@@ 6. Perfect strategies for quantum games
Recall (see [20, Definition 4.1]) that if $X,Y$ are finite sets, a quantum hypergraph over $(X,Y)$ is any subspace $\mathcal{U}\subseteq\overline{\mathbb{C}}^{X}\otimes\mathbb{C}^{Y}$. For a classical hypergraph $E\subseteq X\times Y$, let
@@ 6.1. Perfect strategies for quantum implication games
[…]
@@
(27) $\omega_{\mathfrak{E}}(P,Q)=\sup\limits_{\Phi\in\mathfrak{E}}{\rm Tr}(\Phi(P)Q)$
@@
Suppose that $\Gamma(T)=\sum\limits_{p=1}^{t}N_{p}TN_{p}^{*}$ and $\mathcal{E}(S)=\sum\limits_{r=1}^{s}M_{r}SM_{r}^{*}$. By Lemma 6.2, as $\mathcal{E}$ is a perfect strategy for $\varphi_{P_{1}\rightarrow Q_{1}}$ then $M_{r}\in\tilde{\mathcal{U}}_{P_{1},Q_{1}}$ for each $r=1,\ldots,s$. By construction (see [20]), the Kraus operators of $\Gamma[\mathcal{E}]$ are given by operators $N_{p}[M_{r}]:\mathbb{C}^{X_{2}Y_{2}}\rightarrow\mathbb{C}^{A_{2}B_{2}}$, dependent on $N_{p}$ and $M_{r}$ for each $p=1,\ldots,t,r=1,\ldots,s$. Furthermore, as $\overline{\mathcal{U}}_{P_{1},Q_{1}}\rightarrow\mathcal{U}_{P_{2},Q_{2}}$ via $\Gamma$ by [20, Theorem 5.5] we know that $N_{p}[M_{r}]\in\tilde{\mathcal{U}}_{P_{2},Q_{2}}$ for each $p=1,\ldots,t,r=1,\ldots,s$. This means $\tilde{\mathcal{K}}_{\Gamma[\mathcal{E}]}\subseteq\tilde{\mathcal{U}}_{P_{2},Q_{2}}$; applying Lemma 6.2 once again, we see that $\Gamma[\mathcal{E}]$ is a perfect strategy for $\varphi_{P_{2}\rightarrow Q_{2}}$. Furthermore, by Theorem 5.1 we have that $\Gamma[\mathcal{E}]\in\mathcal{Q}_{\rm t}$, whenever $\Gamma\in\mathcal{Q}_{\rm st}$ with $\mathcal{E}\in\mathcal{Q}_{\rm t}$. ∎
We comment on a special case, which links previous results to a particular class of quantum games of great interest. A rank-one quantum game over $(X,Y,A,B)$ is a pair $(\xi,\gamma)$, where $\xi\in\mathbb{C}^{XY}\otimes\ell_{2}^{R}$ and $\gamma\in\mathbb{C}^{AB}\otimes\ell_{2}^{R}$ are unit vectors; such games were introduced and investigated in [11]. For fixed ${\rm t}\in\{\rm loc,q,qa,qc,ns\}$ the ${\rm t}$-value of the game is given by
$\omega_{\rm t}(\xi,\gamma)=\sup\limits_{\Gamma\in\mathcal{Q}_{\rm t}}{\rm Tr}((\Gamma\otimes{\rm id}_{RR})(\xi\xi^{*})\gamma\gamma^{*}).$
Note that, in the latter definition, we assume $\Gamma:M_{XY}\rightarrow M_{AB}$ as a quantum channel. It was shown (see [11, Proposition 3.1]) that, for rank-one quantum games it is enough to assume that $R$ is a finite set when studying the corresponding game value. Thus, for unit vectors $\xi\in\mathbb{C}^{XYR}$ and $\gamma\in\mathbb{C}^{ABR}$, we may define $\mathcal{U}_{\xi,\gamma}:=\mathcal{U}_{\xi\xi^{*},\gamma\gamma^{*}}$ as before, considering it as a quantum hypergraph over $(XYR,ABR)$. From here, the following result should seem clear in context.
Theorem 6.4.
Let $R,X_{i},Y_{i},A_{i},B_{i},i=1,2$ be finite sets, $\xi_{i}\in\mathbb{C}^{X_{i}Y_{i}}\otimes\mathbb{C}^{R},\gamma_{i}\in\mathbb{C}^{A_{i}B_{i}}\otimes\mathbb{C}^{R},i=1,2$ be unit vectors, and ${\rm t}\in\{\rm loc,q,qa,qc,ns\}$. If $\overline{\mathcal{U}}_{\xi_{1},\gamma_{1}}\rightarrow_{\rm st}\mathcal{U}_{\xi_{2},\gamma_{2}}$ via $\Gamma\otimes{\rm id}_{RR}$, and if $\mathcal{E}\otimes{\rm id}_{R}$ is a perfect strategy for $(\xi_{1},\gamma_{1})$ then $\Gamma[\mathcal{E}]\otimes{\rm id}_{R}$ is a perfect strategy for $(\xi_{2},\gamma_{2})$.
Proof.
The claim follows after arguing as in Theorem 6.3, once we use the fact that $\Gamma\otimes{\rm id}_{RR}\in\mathcal{Q}_{\rm st}$ and $\mathcal{E}\otimes{\rm id}_{R}\in\mathcal{Q}_{\rm t}$ whenever $\Gamma\in\mathcal{Q}_{\rm st},\mathcal{E}\in\mathcal{Q}_{\rm t}$, and that
$(\Gamma\otimes{\rm id}_{RR})[\mathcal{E}\otimes{\rm id}_{R}]=\Gamma[\mathcal{E}]\otimes{\rm id}_{R}.$
Remark 6.5.
We note that the previous result(s) pertain to perfect strategies only. Specifically, if $\omega_{\rm t}(\xi_{1},\gamma_{1})=1$ while $\omega_{\rm t}(\xi_{2},\gamma_{2})<1$ for some ${\rm t}\in\{\rm loc,q,qa,qc,ns\}$ this would automatically imply the non-existence of a quantum ${\rm t}$-homomorphism $\overline{\mathcal{U}}_{\xi_{1},\gamma_{1}}\rightarrow\mathcal{U}_{\xi_{2},\gamma_{2}}$ of the form $\Gamma\otimes{\rm id}_{RR}$; it may or may not be possible to find some ${\rm t}$-homomorphism of the form $\Phi:M_{X_{2}Y_{2}A_{1}B_{1}RR}\rightarrow M_{X_{1}Y_{1}A_{2}B_{2}RR}$, which is not tensored with an ancillary identity operator.
@@ 6.2. Classical-to-quantum non-local games
In this subsection, we consider a way to transfer strategies from classical non-local games to quantum ones, and how a perfect strategy remains perfect under this transfer even when both games are not quantum. For a fixed ${\rm t}\in\{\rm loc,q,qa,qc,ns\}$ to make sense of how we can take a classical strategy $\mathcal{E}\in\mathcal{C}_{\rm t}$ and “apply” it to a legitimately quantum game, the most natural way is to first use [38, Remark 8.1] to consider $\mathcal{E}\in\mathcal{Q}_{\rm t}$, and then use $\Gamma\in\mathcal{Q}_{\rm st}$ as defined in Definition 4.4 to construct a quantum strategy $\Gamma[\mathcal{E}]$ via Theorem 5.1.
Lemma 6.6.
@@
First, assume $\mathcal{E}$ is a perfect strategy for classical non-local game $E$. Viewing $\mathcal{E}$ as a quantum channel, by [20, Remark 4.3]
@@
Assuming the notation and setup as in the proof of Theorem 6.3, by Lemma 6.6 we know that $M_{r}\in\hat{\mathcal{U}}_{E}^{*}$ for each $r=1,\ldots,s$. If the Kraus operators for $\Gamma[\mathcal{E}]$ are given by $N_{p}[M_{r}]$ for $p=1,\ldots,t$ and $r=1,\ldots,s$ then by [20, Theorem 5.5] we know that $N_{p}[M_{r}]\in\tilde{\mathcal{U}}_{P,Q}$ for each $p=1,\ldots,t$ and $r=1,\ldots,s$. By Lemma 6.2, this means $\Gamma[\mathcal{E}]$ is a perfect strategy for $\varphi_{P\rightarrow Q}$. Furthermore, by Theorem 5.1 and Proposition 4.7 we have that $\Gamma[\mathcal{E}]\in\mathcal{Q}_{\rm t}$ whenever $\Gamma\in\mathcal{Q}_{\rm st}$ with $\mathcal{E}\in\mathcal{C}_{\rm t}$. ∎
@@ 6.3. Quantum graph games
(ii) Assume the notation and setup as in the proof of Theorem 6.3; by the aforementioned theorem, we already know that $\Gamma[\mathcal{E}]$ is a perfect strategy for the graph homomorphism game $\mathcal{U}_{2}\rightarrow\mathcal{V}_{2}$. To show that $\Gamma[\mathcal{E}]^{*}$ is a perfect strategy for $\mathcal{V}_{2}\rightarrow\mathcal{U}_{2}$, first note that $\Gamma[\mathcal{E}]^{*}=\Gamma^{*}[\mathcal{E}^{*}]$ (by construction of simulated channels). Furthermore, by definition of $\mathcal{U}_{P_{2},Q_{2}}$ we may easily verify that $\mathcal{U}_{Q_{2},P_{2}}=\overline{\mathcal{U}}_{P_{2},Q_{2}}$. With the use of [20, Lemma 2.1], and as $\theta^{-1}(N_{p}[M_{r}])\in\mathcal{U}_{P_{2},Q_{2}}$, we have that $\theta^{-1}((N_{p}[M_{r}])^{*})=\theta^{-1}(N_{p}^{*}[M_{r}^{*}])\in\overline{\mathcal{U}}_{P_{2},Q_{2}}=\mathcal{U}_{Q_{2},P_{2}},p=1,\ldots,t$ and $r=1,\ldots,s$. Thus, by Lemma 6.2 once more we see $\tilde{\mathcal{K}}_{\Gamma[\mathcal{E}]^{*}}\subseteq\tilde{\mathcal{U}}_{Q_{2},P_{2}}$. This means $\Gamma[\mathcal{E}]^{*}$ is a perfect strategy for $\mathcal{V}_{2}\rightarrow\mathcal{U}_{2}$. That $\Gamma[\mathcal{E}],\Gamma[\mathcal{E}]^{*}$ are both in $\mathcal{Q}_{\rm t}$ follows from Theorem 5.2. Together, these show $\mathcal{U}_{2}\cong_{\rm t}\mathcal{V}_{2}$ via $\Gamma[\mathcal{E}]$, as claimed. ∎
6.4. An algebraic approach
Definition 6.10.
Let $X,Y,A,B$ be finite sets. We let $\mathcal{C}_{XY,AB}$ be the universal ${\rm C}^{*}$-algebra generated by elements $p_{xx^{\prime},yy^{\prime}}^{aa^{\prime},bb^{\prime}},x,x^{\prime}\in X,y,y^{\prime}\in Y,a,a^{\prime}\in A,b,b^{\prime}\in B$ subject to the following relations:
(28) $\sum_{a\in A}p_{xx,yy^{\prime}}^{aa,bb^{\prime}}=\sum_{a\in A}p_{x^{\prime}x^{\prime},yy^{\prime}}^{aa,bb^{\prime}},\;\;\;\;\text{ for \;}x,x^{\prime}\in X$

### v2

@@ 5. Perfect strategies for various quantum games
Recall (see [21, Definition 4.1]) that if $X,Y$ are finite sets, a quantum hypergraph over $(X,Y)$ is any subspace $\mathcal{U}\subseteq\overline{\mathbb{C}}^{X}\otimes\mathbb{C}^{Y}$. For a classical hypergraph $E\subseteq X\times Y$, let
@@ 5.1. Perfect strategies for quantum implication games
$\varphi_{P\rightarrow Q}(\tilde{P})=\begin{cases}Q,\;\;\;\;{\rm if\;}0\neq\tilde{P}\leq P,\\ 0,\;\;\;\;{\rm if\;}\tilde{P}=0,\\ I,\;\;\;\;{\rm otherwise.}\end{cases}$
@@
$\omega_{\mathfrak{E}}(P,Q)=\sup\limits_{\Phi\in\mathfrak{E}}{\rm Tr}(\Phi(P)Q)$ (24)
@@
Suppose that $\Gamma(T)=\sum\limits_{p=1}^{t}N_{p}TN_{p}^{*}$ and $\mathcal{E}(S)=\sum\limits_{r=1}^{s}M_{r}SM_{r}^{*}$. By Lemma 5.2, as $\mathcal{E}$ is a perfect strategy for $\varphi_{P_{1}\rightarrow Q_{1}}$ then $M_{r}\in\tilde{\mathcal{U}}_{P_{1},Q_{1}}$ for each $r=1,\ldots,s$. By construction (see [21]), the Kraus operators of $\Gamma[\mathcal{E}]$ are given by operators $N_{p}[M_{r}]:\mathbb{C}^{X_{2}Y_{2}}\rightarrow\mathbb{C}^{A_{2}B_{2}}$, dependent on $N_{p}$ and $M_{r}$ for each $p=1,\ldots,t,r=1,\ldots,s$. Furthermore, as $\overline{\mathcal{U}}_{P_{1},Q_{1}}\rightarrow\mathcal{U}_{P_{2},Q_{2}}$ via $\Gamma$ by [21, Theorem 5.5] we know that $N_{p}[M_{r}]\in\tilde{\mathcal{U}}_{P_{2},Q_{2}}$ for each $p=1,\ldots,t,r=1,\ldots,s$. This means $\tilde{\mathcal{K}}_{\Gamma[\mathcal{E}]}\subseteq\tilde{\mathcal{U}}_{P_{2},Q_{2}}$; applying Lemma 5.2 once again, we see that $\Gamma[\mathcal{E}]$ is a perfect strategy for $\varphi_{P_{2}\rightarrow Q_{2}}$. Furthermore, by Theorem 4.1 we have that $\Gamma[\mathcal{E}]\in\mathcal{Q}_{\rm t}$, whenever $\Gamma\in\mathcal{Q}_{\rm st}$ with $\mathcal{E}\in\mathcal{Q}_{\rm t}$. ∎
@@ 5.2. Classical-to-quantum non-local games
In this subsection, we consider a way to transfer strategies from classical non-local games to quantum ones, and how a perfect strategy remains perfect under this transfer even when both games are not quantum. For a fixed ${\rm t}\in\{\rm loc,q,qa,qc,ns\}$ to make sense of how we can take a classical strategy $\mathcal{E}\in\mathcal{C}_{\rm t}$ and “apply” it to a legitimately quantum game, the most natural way is to first use [39, Remark 8.1] to consider $\mathcal{E}\in\mathcal{Q}_{\rm t}$, and then use $\Gamma\in\mathcal{Q}_{\rm st}$ as defined in Definition 3.7 to construct a quantum strategy $\Gamma[\mathcal{E}]$ via Theorem 4.1.
Lemma 5.4
@@
First, assume $\mathcal{E}$ is a perfect strategy for classical non-local game $E$. Viewing $\mathcal{E}$ as a quantum channel, by [21, Remark 4.3]
@@
Assuming the notation and setup as in the proof of Theorem 5.3, by Lemma 5.4 we know that $M_{r}\in\hat{\mathcal{U}}_{E}^{*}$ for each $r=1,\ldots,s$. If the Kraus operators for $\Gamma[\mathcal{E}]$ are given by $N_{p}[M_{r}]$ for $p=1,\ldots,t$ and $r=1,\ldots,s$ then by [21, Theorem 5.5] we know that $N_{p}[M_{r}]\in\tilde{\mathcal{U}}_{P,Q}$ for each $p=1,\ldots,t$ and $r=1,\ldots,s$. By Lemma 5.2, this means $\Gamma[\mathcal{E}]$ is a perfect strategy for $\varphi_{P\rightarrow Q}$. Furthermore, by Theorem 4.1 and Proposition 3.10 we have that $\Gamma[\mathcal{E}]\in\mathcal{Q}_{\rm t}$ whenever $\Gamma\in\mathcal{Q}_{\rm st}$ with $\mathcal{E}\in\mathcal{C}_{\rm t}$. ∎
@@ 5.3. Quantum graph games
(ii) Assume the notation and setup as in the proof of Theorem 5.3; by the aforementioned theorem, we already know that $\Gamma[\mathcal{E}]$ is a perfect strategy for the graph homomorphism game $\mathcal{U}_{2}\rightarrow\mathcal{V}_{2}$. To show that $\Gamma[\mathcal{E}]^{*}$ is a perfect strategy for $\mathcal{V}_{2}\rightarrow\mathcal{U}_{2}$, first note that $\Gamma[\mathcal{E}]^{*}=\Gamma^{*}[\mathcal{E}^{*}]$ (by construction of simulated channels). Furthermore, by definition of $\mathcal{U}_{P_{2},Q_{2}}$ we may easily verify that $\mathcal{U}_{Q_{2},P_{2}}=\overline{\mathcal{U}}_{P_{2},Q_{2}}$. With the use of [21, Lemma 2.1], and as $\theta^{-1}(N_{p}[M_{r}])\in\mathcal{U}_{P_{2},Q_{2}}$, we have that $\theta^{-1}((N_{p}[M_{r}])^{*})=\theta^{-1}(N_{p}^{*}[M_{r}^{*}])\in\overline{\mathcal{U}}_{P_{2},Q_{2}}=\mathcal{U}_{Q_{2},P_{2}},p=1,\ldots,t$ and $r=1,\ldots,s$. Thus, by Lemma 5.2 once more we see $\tilde{\mathcal{K}}_{\Gamma[\mathcal{E}]^{*}}\subseteq\tilde{\mathcal{U}}_{Q_{2},P_{2}}$. This means $\Gamma[\mathcal{E}]^{*}$ is a perfect strategy for $\mathcal{V}_{2}\rightarrow\mathcal{U}_{2}$. That $\Gamma[\mathcal{E}],\Gamma[\mathcal{E}]^{*}$ are both in $\mathcal{Q}_{\rm t}$ follows from Theorem 4.2. Together, these show $\mathcal{U}_{2}\cong_{\rm t}\mathcal{V}_{2}$ via $\Gamma[\mathcal{E}]$, as claimed. ∎
6. Characterization of SQNS correlations and applications to concurrent games
In the final section, we wish to link strongly quantum no-signalling correlations to the (multi-variate) tensor product of operator systems. In order to do so, we first briefly recall some facts about operator systems and related constructions. If $\mathcal{S}$ and $\mathcal{T}$ are operator systems, we call $\mathcal{S}$ and $\mathcal{T}$ completely order isomorphic and write $\mathcal{S}\cong_{\rm c.o.i}\mathcal{T}$ if there exists a unital completely positive bijection $\phi:\mathcal{S}\rightarrow\mathcal{T}$ with completely positive inverse. We write $\mathcal{S}\subseteq_{\rm c.o.i}\mathcal{T}$ if $\mathcal{S}\subseteq\mathcal{T}$ and the inclusion map $\iota:\mathcal{S}\hookrightarrow\mathcal{T}$ is a complete order isomorphism onto its range. The three main types of operator system tensor products that will be used in the sequel are given as follows:
  • (i)
  • (ii)
  • (iii)
More details about each tensor product may be found in [27]; the construction of the multivariate tensor product of each type, with explicit descriptions of their matricial cones, may be found in [20, Section 7].
We recall the notion of a coproduct of operator systems: if $\mathcal{S}$ and $\mathcal{T}$ are two operator systems, their coproduct $\mathcal{S}\oplus_{1}\mathcal{T}$ is the unique (up to isomorphism) operator system equipped with complete order embeddings $\iota_{\mathcal{S}}:\mathcal{S}\rightarrow\mathcal{S}\oplus_{1}\mathcal{T}$ and $\iota_{\mathcal{T}}:\mathcal{T}\rightarrow\mathcal{S}\oplus_{1}\mathcal{T}$ which satisfies the following unversal property: For each ucp map $\phi:\mathcal{S}\rightarrow\mathcal{R}$ and $\psi:\mathcal{T}\rightarrow\mathcal{R}$, where $\mathcal{R}$ is an operator system, there exists a unique ucp map $\varphi:\mathcal{S}\oplus_{1}\mathcal{T}\rightarrow\mathcal{R}$ such that $\varphi(\iota_{\mathcal{S}}(s))=\phi(s)$ and $\varphi(\iota_{\mathcal{T}}(t))=\psi(t)$ for every $s\in\mathcal{S},t\in\mathcal{T}$. For more properties of the coproduct of operator systems, we refer the reader to [25, Section 8].
We also will need to recall the universal operator system for stochastic operator matrices, introduced in [39]. A $\it ternary\;ring\;of\;operators\;(TRO)$} is a subspace $\mathcal{V}\subseteq\mathcal{B}(H,K)$ for some Hilbert spaces $H$ and $K$, such that $ST^{*}R\in\mathcal{V}$ whenever $S,T,R\in\mathcal{V}$ (see [3, Section 4.3] or [41]). Let $X$ and $A$ be finite sets, and let $\mathcal{V}_{X,A}$ be the universal TRO generated by the entries $v_{a,x}$ of a block operator isometry $V=(v_{a,x})_{a\in A,x\in X}$. That is, $\mathcal{V}_{X,A}$ is the universal TRO with generators $v_{a,x},a\in A,x\in X$, ternary operator $\cdot,\cdot,\cdot]:\mathcal{V}_{X,A}\times\mathcal{V}_{X,A}\times\mathcal{V}_{X,A}\rightarrow\mathcal{V}_{X,A}$ given by $u,v,w]=uv^{*}w$ for $u,v,w\in\mathcal{V}_{X,A}$, and relations
Let $\mathfrak{C}_{X,A}$ be the unital *-algebra, generated by the set $\{v_{a,x}^{*}v_{a^{\prime},x^{\prime}}:\;x,x^{\prime}\in X,a,a^{\prime}\in A\}$, and $\mathcal{C}_{X,A}$ be the $\it right\;C^{*}-algebra$} of $\mathcal{V}_{X,A}$; up to a *-isomorphism, we may view
We also will consider the space
$\mathcal{S}_{X,A}:={\rm span}\{e_{x,x,a,a}:\;x\in X,a\in A\},$
viewed as an operator subsystem inside $\mathcal{T}_{X,A}$. To help distinguish between operator systems $\mathcal{T}_{X,A}$ and $\mathcal{T}_{Y,B}$, we will denote the canonical generators of $\mathcal{T}_{X,A}$ by $e_{x,x^{\prime},a,a^{\prime}},x,x^{\prime}\in X,a,a^{\prime}\in A$ and $\mathcal{T}_{Y,B}$ by $f_{y,y^{\prime},b,b^{\prime}},y,y^{\prime}\in Y,b,b^{\prime}\in B$.
[…]
be viewed as an operator system in $\mathcal{C}_{X}$. Furthermore, we let
$\mathcal{S}_{X}:={\rm span}\{e_{x,x,a,a}:\;x\in X,a\in A\}$
be viewed as an operator subsystem of $\mathcal{T}_{X}$.
Proposition 6.1
Proof.
Let $K$ be a Hilbert space, $V:H\rightarrow K$ be an isometry, and
be mutually commuting stochastic operator matrices on $K$ satisfying (9). Define linear maps $\phi:\mathcal{T}_{X,A}\rightarrow\mathcal{B}(K)$ (resp. $\psi:\mathcal{T}_{Y,B}\rightarrow\mathcal{B}(K)$) by $\phi(e_{x,x^{\prime},a,a^{\prime}})=E_{xx^{\prime},aa^{\prime}}$ (resp. $\psi(f_{y,y^{\prime},b,b^{\prime}})=F_{yy^{\prime},bb^{\prime}}$). By [39, Theorem 5.2], both $\phi$ and $\psi$ are unital and completely positive. By the definition of the commuting tensor product of operator spaces, the map $\phi\cdot\psi:\mathcal{T}_{X,A}\otimes_{\rm c}\mathcal{T}_{Y,B}\rightarrow\mathcal{B}(K)$ given by $(\phi\cdot\psi)(u\otimes v)=\phi(u)\psi(v)$ is (unital and) completely positive as well. Set
Conversely, suppose that $\gamma:\mathcal{T}_{X,A}\otimes_{\rm c}\mathcal{T}_{Y,B}\rightarrow\mathcal{B}(H)$ is a unital completely positive map. By [39, Corollary 5.3] and [27, Theorem 6.4], $\mathcal{T}_{X,A}\otimes_{\rm c}\mathcal{T}_{Y,B}\subseteq_{\rm c.o.i}\mathcal{C}_{X,A}\otimes_{\rm max}\mathcal{C}_{Y,B}$. Apply Arveson’s Extension Theorem to obtain a completely positive map $\tilde{\gamma}:\mathcal{C}_{X,A}\otimes_{\rm max}\mathcal{C}_{Y,B}\rightarrow\mathcal{B}(H)$ extending $\gamma$. If we then apply Stinespring’s Theorem, we may write
Remark 6.2.
By Remark 3.2 and [20, Remark 5.3], there exist strongly stochastic operator matrices which are not dilatable.
Proposition 6.3
Proof.
Using [7, Theorem 3.4] and [27, Theorem 6.4], argue exactly as in Proposition 6.1 but with bistochastic operator matrices in place of stochastic operator matrices and $\mathcal{T}_{X}$ (resp. $\mathcal{T}_{Y}$) in place of $\mathcal{T}_{X,A}$ (resp. $\mathcal{T}_{Y,B}$). ∎
6.1. Representations of SQNS correlations via operator systems
[…]
${\rm max}-\otimes_{j=1}^{k}\mathcal{S}_{j}:=\mathcal{S}_{1}\otimes_{\rm max}\cdots\otimes_{\rm max}\mathcal{S}_{k}.$
The $k$-fold commuting tensor product was initially defined in [20, Section 7]: if $H$ is a Hilbert space and $\phi_{j}:\mathcal{S}_{j}\rightarrow\mathcal{B}(H)$ are completely positive maps for $j=1,\ldots,k$, we call the family $(\phi_{j})_{j=1}^{k}$ commuting if $\phi_{i}$ and $\phi_{j}$ have mutually commuting ranges whenever $i\neq j$, for $1\leq i,j\leq k$. Given a commuting family $(\phi_{j})_{j=1}^{k}$, we define the linear map $\prod_{j=1}^{k}\phi_{j}:\otimes_{j=1}^{k}\mathcal{S}_{j}\rightarrow\mathcal{B}(H)$ via
$\bigg(\prod_{j=1}^{k}\phi_{j}\bigg)(\otimes_{j=1}^{k}u_{j}):=\prod_{j=1}^{k}\phi_{j}(u_{j}),\;\;\;\;u_{j}\in\mathcal{S}_{j},j\in[k].$
The positive cones for ${\rm c}-\otimes_{j=1}^{k}\mathcal{S}_{j}$ are then determined by all elements $u\in M_{n}(\otimes_{j=1}^{k}\mathcal{S}_{j})$ which are positive in $M_{n}(\mathcal{B}(H))$ under all mutually commuting families and for all choices of Hilbert space $H$.
[…]
Theorem 6.4
Proof.
[…]
Lemma 6.5
Proof.
This proof relies on similar ideas as the proof of [39, Lemma 6.6]; we include the details for the convenience of the reader. We only consider the case when $k=3$; all others will follow similarly. Let $D$ be the set containing all sums of elementary tensors $v_{1}\otimes v_{2}\otimes v_{3}$ with $v_{i}\in V_{i}^{+},\;i=1,2,3$. We claim that if, for every $\epsilon>0$, there exists $u_{\epsilon}\in D$ such that $\|u_{\epsilon}\|\rightarrow 0$ as $\epsilon\rightarrow 0$ and $u+u_{\epsilon}\in D$ for each $\epsilon>0$, then $u\in D$. We may, without loss of generality, further assume that $\|u_{\epsilon}\|\leq 1$ for all $\epsilon>0$. Set $K=2{\rm dim}(V_{1}){\rm dim}(V_{2}){\rm dim}(V_{3})+1$ and, using Carathéodory’s Theorem, start by writing
where $v_{i,j}^{(\epsilon)}\in V_{i}^{+}$ and $\|v_{1,j}^{(\epsilon)}\|=\|v_{2,j}^{(\epsilon)}\|=\|v_{3,j}^{(\epsilon)}\|$ for $i=1,2,3$, $j=1,\ldots,K$ and all $\epsilon>0$. As $v_{1,j}^{(\epsilon)}\otimes v_{2,j}^{(\epsilon)}\otimes v_{3,j}^{(\epsilon)}\leq u+u_{\epsilon}$ and $\|u+u_{\epsilon}\|\leq\|u\|+1$ for all $\epsilon>0$, we get that $\|v_{i,j}^{(\epsilon)}\|\leq\sqrt[3]{\|u\|+1}$ for $i=1,2,3$ and $j=1,\ldots,K$. Using the fact that all of our AOU spaces are finite-dimensional, by compactness we can also assume $v_{i,j}^{(\epsilon)}\rightarrow v_{i,j}$ as $\epsilon\rightarrow 0$ for $i=1,2,3$ and all $j=1,\ldots,K$. We may then conclude that $u=\sum_{j=1}^{K}v_{1,j}\otimes v_{2,j}\otimes v_{3,j}\in D$.
Let
$S_{0}=\sum\limits_{p=1}^{\ell}a_{p}\otimes v_{p}^{(1)},\;\;\;\;T_{0}=\sum\limits_{q=1}^{s}b_{q}\otimes v_{q}^{(2)},\;\;\;\;U_{0}=\sum\limits_{r=1}^{t}c_{r}\otimes v_{r}^{(3)},$ (25)
Now suppose first that $S\in M_{n}({\rm OMAX}(V_{1}))^{+}$ and $\alpha\in M_{1,nmd}$. By the description of the multivariate maximal tensor product as provided in [20, Proposition 7.2], if $\epsilon>0$ then $S+\epsilon 1_{n}$ has the form of $S_{0}$ as in (25). Therefore,
$\alpha(S\otimes T_{0}\otimes U_{0})\alpha^{*}\in D.$
For a linear functional
$\big(s(e_{x_{2},x_{2}^{\prime},x_{1},x_{1}^{\prime}}\otimes e_{y_{2},y_{2}^{\prime},y_{1},y_{1}^{\prime}}\otimes e_{a_{1},a_{1}^{\prime},a_{2},a_{2}^{\prime}}\otimes e_{b_{1},b_{1}^{\prime},b_{2},b_{2}^{\prime}})\big)_{x_{2}x_{2}^{\prime},y_{2}y_{2}^{\prime},a_{1}a_{1}^{\prime},b_{1}b_{1}^{\prime}}^{x_{1}x_{1}^{\prime},y_{1}y_{1}^{\prime},a_{2}a_{2}^{\prime},b_{2}b_{2}^{\prime}}.$ (26)
Theorem 6.6
The map $s\mapsto\Gamma_{s}$ is an affine isomorphism from
  • (i)
  • (ii)
  • (iii)
  • (iv)
Proof.
(i) Let $\Gamma\in\mathcal{Q}_{\rm sns}$. If
for all $x_{2},x_{2}^{\prime}\in X_{2}$. This equality implies